Imperial Devices (ID) has offered to supply the state government with one model of its security screening device at “cost plus 20 percent.” ID operates a manufacturing plant that can produce 66,000 devices per year, but it normally produces 60,000. The costs to produce 60,000 devices follow:

Total Cost Cost per
Device
Production costs:
Materials $ 4,500,000 $ 75
Labor 9,000,000 150
Supplies and other costs that will vary with production 2,700,000 45
Indirect cost that will not vary with production 2,700,000 45
Variable marketing costs 1,800,000 30
Administrative costs (will not vary with production) 5,400,000 90
Totals $ 26,100,000 $ 435

Based on these data, company management expects to receive $522 (= $435 × 120 percent) per monitor for those sold on this contract. After completing 500 monitors, the company sent a bill (invoice) to the government for $261,000 (= 500 monitors × $522 per monitor).

The president of the company received a call from a state auditor, who stated that the per monitor cost should be:

Materials $ 75
Labor 150
Supplies and other costs that will vary with production 45
$ 270

Therefore, the price per monitor should be $324 (= $270 × 120 percent). The state government ignored marketing costs because the contract bypassed the usual selling channels.

Required:

For each of the four situations, calculate the cost basis per device based on the information shown above. (Round intermediate calculations and final answers to 2 decimal places.)

Options:

  1. Only the differential production costs could be considered as the cost basis.
  2. The total cost per device for normal production of 60,000 devices could be used as the cost basis.
  3. The total cost per device for production of 66,000 devices, excluding marketing costs, could be used as the cost basis.
  4. The total cost per device for production of 66,000 devices, including marketing costs, could be used as the cost basis.
Answer #1

60,000 Ansever Page No ③ Production costo 66000 Materials 4500,000 75 X 66000 - 4950,000 Labour 9000,000 150 X 66000 = 9900,0Option o Page No ③ Per device cost basis = 75 + 150 + 45 + 2? + 2400,000 66000 5400,000 66000 = 75 +15+ 45 + 40.91 +81.82 = 3